Fermi National Laboratory

Accelerator Update for December 17, 2007 - December 19, 2007


Monday December 17

The day shift began with Operations monitoring the Tevatron (TeV) store 5808 (111.61E30), the Recycler stash (48.5E10), with the Antiproton Source (Pbar) stacking (40.7E10), with the Main Injector (MI), Booster, and the Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, and with NuMI , SciBooNE / MiniBooNE experiments taking beam.

Operations reported at 8:39 AM that a Pbar vacuum pump (D:IP708) had been tripping off frequently for the past couple of days. Operators reported that there probably was a small vacuum leak in the system.

A Pbar target expert reported at 9:32 AM that he was going to move the target to a new spot to increase the antiproton yield.

Booster experts began tuning their accelerator at 10:01 AM to reduce the BEX bump current.

At 10:34 AM, the MI high level RF system tripped off due to bad vacuum. It would reset, but would not stay on. An RF expert investigated and reported that vacuum bursts were keeping the system off and were being caused by a vacuum pump (IP605D).

At 11 AM, a Pbar expert reported that the D:IP708 vacuum pump was being removed from the vacuum permit system and that the next time it tripped off, Operations should leave it off.

The High Voltage Coordinator reported at 11:12 AM that the Feeder 36 fault was found in a manhole upstream of the Proton PS3 area. This bad splice will be repaired on Tuesday.

At 11:18 AM, two Pbar vacuum pumps began tripping off (IP403 and IP712). Later, a Pbar expert said to leave IP712 off also.

The Neutron Therapy Facility (NTF) reported at 11:32 AM that they finished treating their last patient for the day.

Switchyard experts reported at 11:36 AM that they had completed their studies.

At 1:13 PM, Pbar vacuum got so bad that a Pbar expert asked operators to keep resetting IP708.

A MI expert began a slipstacking study at 2:10 PM.

Vacuum experts reported at 3:50 PM that there was a vacuum leak in the Pbar AP2 line near the IP708 vacuum pump that was causing the pump to trip off. Operators will keep resetting IP708.

Operations reported at 4:03 PM that the Recycler had stashed antiprotons three times during the day shift.

MI studies ended around 5:52 PM.

At 8:27 PM, a Linac quadrupole power supply (L:QPS107) began alarming. Operators swapped out its control card.

At 10:25 PM, operators found several Pbar vacuum pumps off. They all reset.

At 11:24 PM, QPS107 went flakey again. Operators swapped out its control card again.

Operations reported at midnight that the Recycler had stashed antiprotons four times during the evening shift.

Tuesday December 18

The midnight shift began with Operations monitoring store 5808 (38.6E30) and the stash (310.7E10), with Pbar stacking (17.1E10), and with NuMI and Sci /MiniBooNE taking beam.

Operations terminated store 5808 at 3:04 AM.

Operations put the TeV into shot setup at 3:21 AM.

Operations prepared MI-31 for an access into the Pelletron at 5:17 AM.

Operations established store 5809 at 5:32 AM with an initial luminosity of 200.31E30.

At 5:28 AM, a MI kicker (I:KPS1B) tripped off for the first of three times in 3.5 hours. It reset.

At 5:48 AM, a Linac Klystron RF station (KRF6) tripped off. It reset.

At 6:50 AM, a Tevatron Electron Lens (TEL1) tripped off. It reset.

At 7:51 AM, KRF3 tripped off. It reset.

Operations reported at 7:57 AM that the Recycler had stashed two times during the midnight shift.

Operations began preparing Booster for an access at 9:41 AM.

Preaccelerator experts shut down the H- Source, which was the back up source, at 10:18 AM in preparation for work scheduled for Wednesday.

Operations prepared NuMI for an access at 11:26 AM.

At 11:52 AM, a Pbar expert reported that they found the vacuum leak on the IP708 vacuum pump in the AP2 line.

Vacuum technicians reported that they had completed their repair on IP708 at 1:30 PM.

Pbar resumed stacking around 2:12 PM.

At 2:45 PM, BRF15 tripped off and would spark on each attempt to reset. Operators bypassed the station and brought BRF19 on line in its place.

At 2:56 PM, KRF3 tripped off. Linac experts investigated and found a 480-volt wire had shorted to a cooling fan.

Operations reported at 4:03 PM that the Recycler had stashed once during the day shift.

Linac resumed sending beam at 4:38 PM.

NuMI resumed taking beam at 4:41 PM.

Sci/MiniBooNE resumed taking beam at 6:58 PM.

Wednesday December 19

The midnight shift began with Operations monitoring store 5809 (50.31E30) and a stash (123.3E10), with Pbar stacking (136.3E10), with experts working on the Pelletron, and with NuMI and Sci/MiniBooNE taking beam.

At 5:40 AM, Operations reported that due to the large stack size, the Pbar stack began to unstack. Operators adjusted tunes and the ARF2 frequency to compensate.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow TeV experts to conduct an end-of-store study and for MI experts to conduct aperture scans.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/19/2007    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory